3x-7y=-3
2x+6y+-34

Answer by checkley77(12844)

3x-7y=-3 multiply by 2
2x+6y=-34 multiply by -3 & add.
------------------
6x-14y=-6
-6x-18y=102
------------
-32y=96
y=96/-32
y=-3 answer.
3x-7*-3=-3
3x+21=-3
3x=-3-21
3x=-24
x=-24/3
x=-8 answer.
2*-8+6*-3=-34
-16-18=-34
-34=-34